# Aligning Large Language Models by On-Policy Self-Judgment

### v2

@@ Appendix B Qualitative Examples
Prompt what is the name of chris tucker first movie
DPO The name of Chris Tucker’s first movie is “Friday”. It was released in 1995. The movie was written by Tucker and Ice Cube, and also stars Ice Cube and Nia Long. RSO Chris Tucker’s first major film role was in the 1995 comedy film “Friday”. He also had roles in “The Fifth Element” and “Money Train”. His most famous films include “Rush Hour”, “The Best Man”, and “Rush Hour 2”. ReST The name of Chris Tucker’s first movie was “Friday”. RAFT The name of Chris Tucker’s first movie is “Friday”, which was released in 1995. RLHF His first film was "Friday", released in 1995. The film was a great success, and helped launch him to fame. It earned him a Golden Globe nomination for Best Actor in a Musical or Comedy, and received critical acclaim. It is considered a classic modern comedy movie, and has become a cult film. A sequel was released in 2007, and a third film in the franchise is planned for 2017. Self-Judge The first movie that Chris Tucker starred in was the 1995 comedy film “Friday”, directed by F. Gary Gray. The film is a classic urban comedy, centering around the adventures of two friends, played by Tucker and Ice Cube, as they spend the day hanging out in their Los Angeles neighborhood. The film was a breakout hit for Tucker, establishing him as a prominent comedic actor, and was followed by a successful sequel in 1997. Tucker then went on to star in several major Hollywood films, including the hit comedy films “Boomerang” and “Tommy” as well as the drama “Rush Hour” and its sequels. He has also had a successful career as a stand-up comedian, performing sold-out shows worldwide.
@@
Prompt Take a look at the contract and extract the parties of the agreement from it. As of the 13th day of December 2021, this Music Recording Agreement ("Agreement") is made between Good Kid, a Toronto-based musical group ("Artist"), and Universal Music Group, a record label with license number 545345 ("Record Label"). The Artist and the Recording Label may be referred to in this Agreement individually as "Parties" and collectively as "Parties." Work under this Agreement will begin on March 15, 2022, and will last for the duration of the Agreement.
JM ($\mathbf{N}$ = 1) The parties of the Music Recording Agreement are: 1. Good Kid, a Toronto-based musical group ("Artist") 2. Universal Music Group, a record label with license number 545345 ("Record Label") Both parties are referred to individually as "Parties" in the Agreement and collectively as "Parties." This means that the Artist and the Record Label are the two main participants in the agreement. JM-PR ($\mathbf{N}$ = 1) The parties involved in the Music Recording Agreement are: * Good Kid, a Toronto-based musical group, represented by the artist. * Universal Music Group, a record label with license number 545345, represented by the record label.) This agreement is effective as of the 13th day of December 2021 and will last until the expiration of the agreement. The work under this agreement will begin on March 15, 2022, and will continue for the duration of the agreement.
JM ($\mathbf{N}$ = 16) Based on the provided Music Recording Agreement, the parties involved in the agreement are: 1. Good Kid, a Toronto-based musical group (Artist). 2. Universal Music Group, a record label with license number 545345 (Record Label). The Agreement refers to the Artist and the Record Label as "Parties" and collectively as "Parties" throughout the document. Therefore, the parties to the Music Recording Agreement are: * Good Kid, a Toronto-based musical group (Artist) * Universal Music Group, a record label with license number 545345 (Record Label) The Agreement outlines the terms and conditions of the recording agreement between the Artist and the Record Label, including the commencement date, the duration of the agreement, and the scope of work. JM-PR ($\mathbf{N}$ = 16) Based on the provided contract, the parties involved in the Music Recording Agreement are: 1. Good Kid, a Toronto-based musical group ("Artist") 2. Universal Music Group, a record label with license number 545345 ("Record Label") These two parties are involved in the agreement and are referred to individually as "Parties" and collectively as "Parties" throughout the document.

### v3

@@ Appendix B Qualitative Examples
Prompt
what is the name of chris tucker first movie
DPO
The name of Chris Tucker’s first movie is “Friday”. It was released in 1995. The movie was written by Tucker and Ice Cube, and also stars Ice Cube and Nia Long.
RSO
Chris Tucker’s first major film role was in the 1995 comedy film “Friday”. He also had roles in “The Fifth Element” and “Money Train”. His most famous films include “Rush Hour”, “The Best Man”, and “Rush Hour 2”.
ReST
The name of Chris Tucker’s first movie was “Friday”.
RAFT
The name of Chris Tucker’s first movie is “Friday”, which was released in 1995.
RLHF
His first film was "Friday", released in 1995. The film was a great success, and helped launch him to fame. It earned him a Golden Globe nomination for Best Actor in a Musical or Comedy, and received critical acclaim. It is considered a classic modern comedy movie, and has become a cult film. A sequel was released in 2007, and a third film in the franchise is planned for 2017.
Self-Judge
The first movie that Chris Tucker starred in was the 1995 comedy film “Friday”, directed by F. Gary Gray. The film is a classic urban comedy, centering around the adventures of two friends, played by Tucker and Ice Cube, as they spend the day hanging out in their Los Angeles neighborhood. The film was a breakout hit for Tucker, establishing him as a prominent comedic actor, and was followed by a successful sequel in 1997. Tucker then went on to star in several major Hollywood films, including the hit comedy films “Boomerang” and “Tommy” as well as the drama “Rush Hour” and its sequels. He has also had a successful career as a stand-up comedian, performing sold-out shows worldwide.
@@
Prompt
Take a look at the contract and extract the parties of the agreement from it.
As of the 13th day of December 2021, this Music Recording Agreement ("Agreement") is made between Good Kid, a Toronto-based musical group ("Artist"), and Universal Music Group, a record label with license number 545345 ("Record Label"). The Artist and the Recording Label may be referred to in this Agreement individually as "Parties" and collectively as "Parties." Work under this Agreement will begin on March 15, 2022, and will last for the duration of the Agreement.
JM ($\mathbf{N}$ = 1)
The parties of the Music Recording Agreement are:
1. Good Kid, a Toronto-based musical group ("Artist")
2. Universal Music Group, a record label with license number 545345 ("Record Label")
Both parties are referred to individually as "Parties" in the Agreement and collectively as "Parties." This means that the Artist and the Record Label are the two main participants in the agreement.
JM-PR ($\mathbf{N}$ = 1)
The parties involved in the Music Recording Agreement are:
* Good Kid, a Toronto-based musical group, represented by the artist.
* Universal Music Group, a record label with license number 545345, represented by the record label.)
This agreement is effective as of the 13th day of December 2021 and will last until the expiration of the agreement. The work under this agreement will begin on March 15, 2022, and will continue for the duration of the agreement.
JM ($\mathbf{N}$ = 16)
Based on the provided Music Recording Agreement, the parties involved in the agreement are:
1. Good Kid, a Toronto-based musical group (Artist).
2. Universal Music Group, a record label with license number 545345 (Record Label).
The Agreement refers to the Artist and the Record Label as "Parties" and collectively as "Parties" throughout the document. Therefore, the parties to the Music Recording Agreement are:
* Good Kid, a Toronto-based musical group (Artist)
* Universal Music Group, a record label with license number 545345 (Record Label)
The Agreement outlines the terms and conditions of the recording agreement between the Artist and the Record Label, including the commencement date, the duration of the agreement, and the scope of work.
JM-PR ($\mathbf{N}$ = 16)
Based on the provided contract, the parties involved in the Music Recording Agreement are:
1. Good Kid, a Toronto-based musical group ("Artist")
2. Universal Music Group, a record label with license number 545345 ("Record Label")
These two parties are involved in the agreement and are referred to individually as "Parties" and collectively as "Parties" throughout the document.